If a1< ...... <an . Find the minimum value of f(x) = (x - ai)² summation from i=1 to n


1
Expert's answer
2021-03-05T07:25:04-0500


Suppose "a_1<a_2<\u22ef<a_n" . Then on each interval "(a_k,a_{k+1}),"


"f(x)=\u2211_{i=1}^k(a_i\u2212x)+\u2211_{i=k+1}^n(x\u2212a_i)=(n\u22122k)x+\u2211_{i=1}^kai\u2212\u2211_{i=k+1}^na_i"


Thus, f(x) is decreasing on "(a_k,a_{k+1})" as long as "n\u22122k>0," i.e. "k<\dfrac{n}{2}" , increasing as soon as "k>\dfrac{n}{2}."


We see there are two cases:


If n is odd; n=2p+1, there is a unique minimum, attained at the middle point "a_{p+1}" , and its value is:


"f(a_{n+1})=\u2212a_1\u2212\u22ef\u2212a_n+a_{n+2}+\u22ef+a_{2p+1}."


If n is even: n=2p, the function attains its minimum on the middle interval "[a_p,a_{p+1}]" and its value is:

"\u2212a_1\u2212\u22ef\u2212a_p+a_{p+1}+\u22ef+a_{2p}."